Skip to content
CalcGospel 國際數學圖譜
返回

CIE 9709 2023 November Paper 62 Q7

A Level / CIE / S2

CIE 9709 2023 Nov Paper 62 Paper · Question 7

题目

Problem

A random variable XX has the distribution Po(2.4)\mathrm{Po}(2.4).

(a) Find P(2≤X<4)\mathrm{P}(2\leq X<4).

(2)

(b) Two independent values of XX are chosen.

Find the probability that both of these values are greater than 11.

(3)

(c) It is given that P(X=r)<P(X=r+1)\mathrm{P}(X=r)<\mathrm{P}(X=r+1).

(i) Find the set of possible values of rr.

(3)

(ii) Hence find the value of rr for which P(X=r)\mathrm{P}(X=r) is greatest.

(1)
题目中文翻译

随机变量 XX 服从 Po(2.4)\mathrm{Po}(2.4) 分布。

(a) 求 P(2≤X<4)\mathrm{P}(2\leq X<4)。

(2)

(b) 独立选取 XX 的两个值。

求这两个值都大于 11 的概率。

(3)

(c) 已知 P(X=r)<P(X=r+1)\mathrm{P}(X=r)<\mathrm{P}(X=r+1)。

(i) 求 rr 的所有可能取值所构成的集合。

(3)

(ii) 由此求使 P(X=r)\mathrm{P}(X=r) 最大的 rr 的值。

(1)
原卷通用作答要求

Answer all questions. You should use a calculator where appropriate.

You must show all necessary working clearly; no marks will be given for unsupported answers from a calculator.

Give non-exact numerical answers correct to 3 significant figures, or 1 decimal place for angles in degrees, unless a different level of accuracy is specified in the question.

请回答所有题目。适当时应使用计算器。必须清楚展示所有必要的计算过程;仅给出计算器答案而没有过程,不予评分。除非题目另有精度要求,非精确数值答案保留 3 位有效数字,以度为单位的角度保留 1 位小数。

解答

(a)

解法一

思路

展开

区间只包含2和3,累加这两个泊松概率。

答题过程

展开 P(2≤X<4)= e−2.4(2.422!+2.433!)= 0.470282…= 0.470(3 s.f.).\begin{align*} \mathrm{P}(2\leq X<4)=&\,e^{-2.4}\bigg(\frac{2.4^2}{2!}+\frac{2.4^3}{3!}\bigg)\\[4mm] =&\,0.470282\ldots\\[4mm] =&\,\boxed{0.470}\quad\text{(3 s.f.).} \end{align*}

(b)

解法一

思路

展开

单个值大于1的概率先用补集求出,两个独立值同时满足的概率为其平方。

答题过程

展开

For one value,

P(X>1)=1−e−2.4(1+2.4).\mathrm{P}(X>1)=1-e^{-2.4}(1+2.4).

By independence, the required probability is

 [1−e−2.4(1+2.4)]2= 0.478254…≈ 0.478(3 s.f.).\begin{align*} &\,\bigg[1-e^{-2.4}(1+2.4)\bigg]^2\\[4mm] =&\,0.478254\ldots\\[4mm] \approx&\,\boxed{0.478}\quad\text{(3 s.f.).} \end{align*}

(c)(i)

解法一

思路

展开

比较相邻泊松概率并约去公共正因子,化成r加1小于2.4;r须为非负整数。

答题过程

展开

For a non-negative integer rr,

P(X=r)= e−2.42.4rr!,P(X=r+1)= e−2.42.4r+1(r+1)!.\begin{align*} \mathrm{P}(X=r)=&\,e^{-2.4}\frac{2.4^r}{r!},\\[4mm] \mathrm{P}(X=r+1)=&\,e^{-2.4}\frac{2.4^{r+1}}{(r+1)!}. \end{align*}

Thus

 P(X=r)<P(X=r+1)⟺ 1<2.4r+1⟺ r<1.4.\begin{align*} &\,\mathrm{P}(X=r)<\mathrm{P}(X=r+1)\\[4mm] \Longleftrightarrow&\,1<\frac{2.4}{r+1}\\[4mm] \Longleftrightarrow&\,r<1.4. \end{align*}

The possible values are therefore r∈{0,1}\boxed{r\in\{0,1\}}.

(c)(ii)

解法一

思路

展开

概率从0到1再到2递增,而从2到3起递减,因此最大值对应2。

答题过程

展开

From part (i), the probabilities increase from r=0r=0 to r=1r=1 to r=2r=2. For every r≥2r\geq2,

P(X=r+1)P(X=r)=2.4r+1<1,\frac{\mathrm{P}(X=r+1)}{\mathrm{P}(X=r)} =\frac{2.4}{r+1}<1,

so they then decrease. Hence P(X=r)\mathrm{P}(X=r) is greatest at r=2\boxed{r=2}.